if temperature of the ideal black bbody increased by 10% the the percentage increase in the quantity of the radiation emitted from its surface will be

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
T2 = (11/10)T1 
Now 
E is proportional to T4
E(2) - E(1) = (11/10)4 (T1)4 - (T1)4
Hence percentage increase = (E2 - E1)/E1
= [(11/10)4 -1] *100
= [(1.1)4 -1] *100
 = [1.464 - 1] * 100
= 0.464 *100
= 46.4 %
 
Regards
Arun (askIITians forum expert)